HUBBLE SPACE TELESCOPE DAILY REPORT #1301 PERIOD COVERED: 0000Z Daily Status Report as of 009/0000Z 1.0 ACTIVITIES SCHEDULED AND ACCOMPLISHED: 1.1 Completed WF/PC-2 5369 (Medium Deep Survey: Cycle 4 Medium High Galactic Latitude) The WF/PC-2 made a series of observations of high latitude objects as part of a medium deep sky survey. The observations were taken in parallel with other prime science observations. The observations were completed as scheduled, and no anomalies were noted. 1.2 Completed WF/PC-2 5778 (Cycle 4 Lyman Alpha Throughput Verification) The WF/PC-2 was used to make ultraviolet observations of standard star BD+75D325 through filter F130LP as part of a Lyman Alpha throughput verification test. The observations were completed as scheduled, and no anomalies were reported. 1.3 Completed WF/PC-2 5354 (Imaging in the Crab Nebula) The WF/PC-2 was used to make images of NGC-1952-BR (aka, M-1, the Crab Nebula) to study ionization structure and to look for hot spots. The observations were completed as planned, and no problems were reported. 1.4 Completed WF/PC-2 5560 (Cycle 4 Calibration: Internal Monitor 1) The WF/PC-2 was used to make several internal bias, K-Spot, and flat calibration observations to monitor the CCD camera electronics, the optical alignment, and to check for contamination. The observations were completed as planned, and no anomalies were noted. 1.5 Completed FGS 5135 (Extragalactic Astrometry and Astrophysics) The FGSs were used to make astrometric observations to obtain data needed to measure the HIPPARCOS reference frame with respect to the extragalactic reference frame. The observations were executed as scheduled, and no anomalies were noted. 1.6 Completed WF/PC-2 5091 (GTO High Latitude Parallel Program - Cycle 4) The WF/PC-2 was used in parallel with the prime science observations to make high galactic latitude observations greater than +/-30 degrees. The observations were executed as scheduled, and no problems were reported. 1.7 Completed WF/PC-2 5446 (The Nuclei of Nearby SO and Spiral Galaxies: A PC Snapshot Imaging Survey) The WF/PC-2 was used to make observations of the core region of galaxy NGC-1527 to provide statistics on the nuclear cusps, stellar condensations, and inner stellar disks and dust lanes. The observations were executed as scheduled, and no anomalies were noted. 1.8 Completed FOS/BL 5711 (Black Hole Binaries in the LMC) The Faint Object Spectrograph (Blue Side) was used to study the LMC binary CAL 87, which is the known eclipsing black hole candidate. The changes in the continuum will be compared with model predictions to learn more about the accretion disk. The proposal completed with no reported anomalies. 1.9 Completed WF/PC-2 5523 (Post-COSTAR FOC/96 Fine Aperture Location) The WF/PC-2 was used to refine the location of the f/96 aperture in the V2-V3 coordinate frame. There were no reported problems during this observation. 1.10 Completed WF/PC-2 5092 (GTO Low Latitude Parallel Program - Cycle 4) The WF/PC-2 was used to make observations taken in parallel with the prime science observations. This is part of a survey to take pictures of a large number of fields within 30 degrees of the galactic plane. The observations were completed as planned, and no problems were reported. A second set of planned observations were lost because of the Safemode entry due to SADE Failure Flag Test at 008/0820Z. (See Note 2.4) 1.11 Completed WF/PC-2 5629 (Cycle 4 Photometric Calibration Monitor 1C: 4-Chip Standard in F170W) The WF/PC-2 was used to make photometric calibration observations of the ultraviolet standard star GRW+70D582. The observations were executed as scheduled, and no anomalies were noted. 1.12 Completed WF/PC-2 5764 (Cycle 4 Calibration: Internal Flats - Repeat) The WF/PC-2 was used to make internal flat calibration images using the VIS lamp and internal sources. The observations were executed as scheduled, and no anomalies were noted. A second set of planned observations were lost because of the Safemode entry due to SADE Failure Flag Test at 008/0820Z. (See Note 2.4) 1.13 Completed FOS/RD 5753 (SINS: The Supernova Intensive Study) This proposal is part of a continuing project to study the development of supernovae and their circumstellar matter. In this case, LMC-SN1987 was observed. The observation completed normally, and there were no reported problems. 1.14 Completed WF/PC-2 5584 (Parallel High Resolution Imaging of Diffuse Objects in the Magellanic Cloud) The WF/PC-2 was used to observe diffuse object LMC-30DOR. Of particular interest were H II regions, planetary nebulae, snares, supershells, and proto-stellar and maser regions. The observations were not completed as planned. *** Safemode entry due to SADE Failure Flag Test at 008/0820Z. (See Note 2.4) 2.0 FLIGHT OPERATIONS SUMMARY: 2.1 Guide Star Acquisitions: Scheduled Acquisitions : 6 Successful : 6 Scheduled Reacquisitions : 10 Successful : 10 2.2 FHST Updates: Scheduled : 17 Successful : 17 2.3 Significant HSTARs: HSTAR 5084: Safemode entry due to SADE Failure Flag Test at 008/082054Z. (See Note 2.4) Science recovery SMS will start at 010/1200Z. 2.4 Operations Notes: The HST entered software sun point safemode at 008/082054Z due to the failure of the SADE torque failure test. The SADE 1 R unit was switched off and SADE 2 turned on by the flight software when the problem was experienced. A health and safety load was uplinked at 008/1744Z, and was for the interval 008/1832Z - 010/0030Z. Another health and safety load has been prepared for the interval 010/0030Z - 010/1200Z. A revised SMS that incorporates certain solar position restrictions has been generated and is to begin execution on Tuesday, Jan. 10 at 7:00 A.M. local. These restrictions are temporary safety measures to enable return to science operations while analysis of the SADE 1 R failure continues. The spacecraft onboard safing configurations have been modified to use the SADE 2 in the event of another safemode entry. Safemode Recovery Status meetings and telecons with ESA are being held today (Monday). Apart from the safing event, ESTR #2 experienced further bit slips. A re-dump produced bit slips, but at different locations. 3.0 ANOMALY STATUS: 3.1 Significant New Anomalies: None 3.2 Significant Unresolved Anomalies: None 4.0 SIGNIFICANT FORTHCOMING EVENTS: HST Science recovery SMS will start at 010/1200Z. /DMH -END-